OpenStudy (anonymous):

solve 6+(2x-7)^2/3=31 (two answers)

OpenStudy (anonymous):

ok are you sure it is to the power or 2/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok \[(2x-7)^{\frac{2}{3}}=25\] \[2x-7=25^{\frac{3}{2}}\] \[2x-7=125\] \[2x=118\] \[x=59\]

OpenStudy (anonymous):

ops \[2x

OpenStudy (anonymous):

\[2x-7=125\] \[2x=132\] \[x=66\]

OpenStudy (anonymous):

sorry. other one \[2x-7=-125\] \[2x=-118\] \[x=-59\]

OpenStudy (anonymous):

okay thats right